# The simple reason why classical gravity can entangle

### v1

@@ Abstract
However, attempts to apply this theory-independent argument to the GIE experiment has been creating controversy ever since it has been proposed [17, 18, 19, 20, 21, 22]. The more recent round of this controversy is the paper by Aziz and Howl [23] followed by the rebuttals by Marletto and Vedral [24] and Diośi [25]. Besides, it is a simple fact that GIE detection is not enough to rule out theories with classical gravity: Trillo and Navascués have shown that even the Diośi-Penrose model [26, 27]—the paradigmatic model of classical-gravity-induced collapse modification of quantum theory—predicts GIE [28].
@@
Wheeler-Feynman absorber theory [29, 30] is an example of a relativistically local theory with no mediator [21]. Even standard relativistically-local QFTs only feature mediation in specific gauges [^2] Take QED, for example: In the Coulomb gauge, also known as radiation gauge, there is a direct interparticle interaction via the (quantum) Coulomb potential which spoils the mediation assumption. [21, 31] and, even then, they do so only approximately [31].
@@ Acknowledgements.
I wish to thank the organisers and participants of the workshop “A look at the interface between gravity and quantum theory 2025,” as well as Anne-Catherine de la Hamette, for useful and encouraging discussions. This project was funded within the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733, and from the Austrian Science Fund (FWF), projects I-6004 and ESP2889224 as well as Grant No. I 5384.

### v2

@@ Abstract
However, attempts to apply this theory-independent argument to the GIE experiment has been creating controversy ever since it has been proposed [17, 18, 19, 20, 21, 22, 23]. The more recent round of this controversy is the paper by Aziz and Howl [24] followed by the rebuttals by Marletto and Vedral [25] and Diośi [26]. Besides, it is a simple fact that GIE detection is not enough to rule out theories with classical gravity: Trillo and Navascués have shown that even the Diośi-Penrose model [27, 28]—the paradigmatic model of classical-gravity-induced collapse modification of quantum theory—predicts GIE [29].
@@
Wheeler-Feynman absorber theory [30, 31] is an example of a relativistically local theory with no mediator [21]. Even standard relativistically-local QFTs only feature mediation in specific gauges [^2] Take QED, for example: In the Coulomb gauge, also known as radiation gauge, there is a direct interparticle interaction via the (quantum) Coulomb potential which spoils the mediation assumption. [21, 32, 33] and, even then, they do so only approximately [32, 33].
@@ Acknowledgements.
I am grateful to Marios Christodoulou, Thomas Galley, Časlav Brukner, Carlo Rovelli, Markus Aspelmeyer, and Richard Howl for many discussions about GIE and the LOCC-like no-go theorems. I also wish to thank the organisers and participants of the workshop “A look at the interface between gravity and quantum theory 2025,” as well as Anne-Catherine de la Hamette, for useful and encouraging discussions. This project was funded within the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733, and from the Austrian Science Fund (FWF), projects I-6004 and ESP2889224 as well as Grant No. I 5384.